# Dynamic optical coherence tomography algorithm for label-free assessment of swiftness and occupancy of intratissue moving scatterers Supplementary Material

### v2

@@ 5.2 Ambiguous LIV curve and unreliable aLIV and Swiftness values
There are two types of LIV curves that inevitably lead to unreliable parameter fitting, and hence produce unreliable aLIV and Swiftness values.
@@ 5.5 Numerical investigation for more complicated scatterer dynamics
In our numerical simulations, we assumed that active transport and cyclosis within the cells are the dominant intratissue activities, and modeled these activities using the random ballistic motion model (Section 3.1.1). However, to comprehensively examine all intratissue activities, a more complex model is necessary. It has been known that various types of intratissue activities occur on different time scales [51]. Consequently, the random ballistic motion model may not represent all types of intratissue activities.
Feng et al.expanded the modeling of intratissue scatterer motions by introducing mono-directional displacement (flow) and diffusion (Brownian) motion models, in addition to the random ballistic motion model[39]. This comprehensive set of motion models may enable the DOCT metrics to be interpreted for a wide spectrum of sample types beyond cellular samples, such as those exhibiting cellular migration and flow, and more aqueous samples.
@@ 6 Conclusion
In this paper, we proposed a new DOCT algorithm. By leveraging the -size dependency of LIV, two DOCT metrics, named aLIV and Swiftness, were derived. Numerical simulations revealed that aLIV is sensitive to the occupancy of dynamic scatterers (dynamic scatterer ratio), whereas Swiftness is sensitive to the speed of the dynamic scatterers. Experimental validations using in vitro tumor spheroids and an ex vivo mouse kidney demonstrated that aLIV and Swiftness images provide more direct insights into the intratissue activities than our conventional DOCT metrics.
Bolstered by numerical-simulation-based interpretation, the multiple DOCT metrics will enhance the utility and applications of DOCT imaging.
@@
Morishita, Mukherjee, El-Sadek, Makita, Yasuno: Sky Technology(F), Nikon(F), Kao Corp.(F), Topcon(F), Panasonic(F), Santec (F). Seesan, Mori, Furukawa, Matsusaka, Fukuda, Lukmanto: None.

### v3

@@ 5.2 Ambiguous LIV curve and unreliable aLIV and Swiftness values
The fitting accuracy of the LIV curve was investigated by calculating the coefficient of determination (R2 value) for the without-drug spheroid data as shown in Fig. 9(a). R2 ranges from 0 for the worst fit to 1 for the best. There are some pixels with small R2 values. For example, the LIV curve with the smallest R2 (R2 = 0.18) (at the location indicated by the arrow) cannot be properly represented by the saturation function [Fig. 9(b)]. Specifically, the LIV is nearly saturated at the first time-window and exhibits a slight subsequent increase. Since the LIV values on this curve is very small, we suspect that this slight subsequent increase is mainly caused by noise rather than intratissue scatterer motions. Although a few pixels show low R2 values, a good mean R2 value was observed within the tissue as 0.90 0.099 (mean standard deviation). This indicates that most LIV curve fittings were performed with reasonable fitting accuracy.
However, the good fit (i.e., large R2) does not always ensure the good stability of the fitting parameters (i.e., saturation level and time constant). There are two types of LIV curves that inevitably lead to unreliable parameter fitting, and hence produce unreliable aLIV and Swiftness values.
@@
Furthermore, for observation purposes, we can mitigate this issue by introducing a new pseudo-color-image generation method which is presented in Ref. [53]. This method assigns Swiftness, aLIV, and OCT intensity to the hue, saturation, and brightness channels, respectively as exemplified in Fig. 9(c), where the image was reprinted from Ref. [53] and is generated from the same dataset with Fig. 9(a).
@@ 5.5 Numerical investigation for more complicated scatterer dynamics
In our numerical simulations, we simplified intracellular and intratissue motions using a single motion model, i.e., random ballistic motion model (Section 3.1.1), although a diffusion (Brownian) motion model can also be considered as the motion model. A previous study using a numerical simulation revealed that LIV and OCDS exhibit similar characteristics under both the random ballistic and diffusion models [41]. Since aLIV and Swiftness are computed based on LIV calculations, we considered that aLIV and Swiftness would show similar results among two motion models if LIV shows similar results. Therefore, we considered that the fundamental characteristics of DOCT contrasts can be clarified from the results of one of the motion models, namely the random ballistic motion model.
However, neither the random ballistic motion model nor the simple diffusion model may represent all types of intratissue activities. To comprehensively examine all intratissue activities, a more complex model might be necessary.
Feng et al.expanded the modeling of intratissue scatterer motions by introducing mono-directional displacement (flow) and diffusion (Brownian) motion models, in addition to the random ballistic motion model [41]. Besides these models, motion models must also account for more ordered or directed intracellular motions, such as cytoplasmic streaming. However, our current random ballistic motion model is expected to recapitulate such motion only as long as the OCT resolution is larger than the domain size of the motion (such as the cell size). Specifically, if the resolution volume encompasses multiple cells, and the directional motions within those cells are not correlated to each other, these ordered or directed motions can be considered random motion to some degree.
In any case, a greater variety of motion models is anticipated. This comprehensive set of motion models could enable the DOCT metrics to be interpreted for a wide spectrum of sample types beyond cellular samples, such as those exhibiting cellular migration and flow, and more aqueous samples.
@@ 6 Conclusion
In this paper, we proposed a new DOCT algorithm. By leveraging the -size dependency of LIV, two DOCT metrics, named aLIV and Swiftness, were derived. Numerical simulations revealed that aLIV is sensitive to the occupancy of dynamic scatterers (dynamic scatterer ratio), whereas Swiftness is sensitive to the speed of the dynamic scatterers. Each metric has a one-to-one relationship with either the occupancy or the speed of moving scatterers, respectively. This indicates that the unique values of occupancy and speed can be determined directly from aLIV and Swiftness. In addition, we emphasize that these two metrics, which are sensitive to different aspects of sample activity, can be obtained simultaneously.
Experimental validations using in vitro tumor spheroids and an ex vivo mouse kidney demonstrated that aLIV and Swiftness images provide more direct insights into the intratissue activities than our conventional DOCT metrics. Furthermore, the experimental results suggest that aLIV and Swiftness are relatively insensitive to the acquisition time window in comparison to our conventional DOCT algorithms (LIV and OCDS). This characteristic might be particularly advantageous for evaluating biological samples requiring short measurement times, such as in vivo samples.
It should be noted that our current numerical simulation uses a simplified model for intratissue scatterer motions, which includes several assumptions. Nevertheless, bolstered by numerical-simulation-based interpretation, the multiple DOCT metrics will enhance the utility and applications of DOCT imaging.
@@
Morishita, Mukherjee, El-Sadek, Makita, Yasuno: Nidek(F), Sky Technology(F), Panasonic(F), Nikon(F), Santec(F), Kao Corp.(F), Topcon(F). Seesan, Mori, Furukawa, Matsusaka, Fukuda, Lukmanto: None.